OpenStudy (tennistar):

What is the simplified form of the expression the square root of the quantity 16 times c to the eighth power. - the square root of c squared. + 2the square root of c squared. + the square root of the quantity 4 times c squared.

OpenStudy (anonymous):

\[\sqrt{16c^8}-\sqrt{c^2}+2\sqrt{c^2}+\sqrt{4c^2}\]

OpenStudy (tennistar):

\[\sqrt{16c ^{8}}-\sqrt{c ^{2}}+2\sqrt{c ^{2}}+\sqrt{4c ^{2}}\]

OpenStudy (tennistar):

yes

OpenStudy (anonymous):

So your having trouble simplifying?

OpenStudy (tennistar):

yes, I know the first part is square and would be\[4c ^{4}\]

OpenStudy (tennistar):

and that I can combine \[\sqrt{c ^{2}}\]

OpenStudy (tennistar):

then I am confused

OpenStudy (anonymous):

\[\sqrt{c^2}=c\]

OpenStudy (anonymous):

You will be left with 2 separate terms

OpenStudy (tennistar):

Ok, I think I get ti now...

OpenStudy (anonymous):

Great. I can check your answer if you like.